OpenStudy (anonymous):

Can anybody help me find the vertex for y=2x^2-4x? Please explain answer.

OpenStudy (anonymous):

Complete the square to put it in vertex form. Easy peasey. Or you can find your 0's (since this one has them) and the vertex will be half-way between them.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

I have a formula for it.

OpenStudy (anonymous):

Clouser, are you familiar with completing the square?

OpenStudy (anonymous):

Which part do you not understand?

OpenStudy (anonymous):

completing the square or finding the 0's?

OpenStudy (anonymous):

So You make it y=2x(x-4)

OpenStudy (anonymous):

Or sorry y=2x(x-2)

OpenStudy (anonymous):

No take 2x^2-4x Factor out the 2. \[y=2(x^2-2x)\] \[y=2(x^2-2x+1-1)\] \[y=2(x-1)^2-2\] So the vertex is (1,-2)

myininaya (myininaya):

vertex form is \[y=a(x-h)^2+k\] (h,k) is vertex

OpenStudy (anonymous):

He was finding the 0s which is ok too

OpenStudy (anonymous):

No use myiniaya's information to read it off my algebra.

OpenStudy (anonymous):

Since this one has them

myininaya (myininaya):

lol

OpenStudy (anonymous):

What do you mean by finding the 0's?

OpenStudy (anonymous):

The two 0's you have there Clouse are 0 and 2. So the vertex will be at x is halfway between those.

OpenStudy (anonymous):

what is wrong with normal old white bread and butter \[-\frac{b}{2a}\]?

OpenStudy (anonymous):

y = 2x(x-2) means that y = 0 when x is 0 or when x is 2

myininaya (myininaya):

\[\frac{0+2}{2}\]

OpenStudy (anonymous):

-b/2a is too hard to try to remember ;p

myininaya (myininaya):

yeah satellite

myininaya (myininaya):

its too hard to remember

OpenStudy (anonymous):

good lord. just use \[-\frac{b}{2a}=\frac{4}{4}=1\]

myininaya (myininaya):

lol

OpenStudy (anonymous):

No! You can't make me! ;p

OpenStudy (anonymous):

Formula for h is -b/2a and formula for k is c-h^2a. :)

OpenStudy (anonymous):

My brain has only enough room for 1 forumla involving quadratic polynomials. So I am using it for the \[x=\frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] one.

OpenStudy (anonymous):

yes i know it is very hard to remember. much harder than completing the square . it has the following hard symbols \[-\] \[b\] / \[2\] \[a\]

OpenStudy (anonymous):

Everything else I derive.

OpenStudy (anonymous):

LOL, thanks guys. Just discovered this site today. It is like having lots of tutors.

OpenStudy (anonymous):

Sorry for the tangent Clouser.

OpenStudy (anonymous):

very difficult. must easier to factor and keep track of what you added and subtracted to complete the square. or take the derivative, set equal zero and solve, or find the point between the two roots. but it is still \[-\frac{b}{2a}\]

OpenStudy (anonymous):

Satellite, from this equation how do you find the axis of symmetry?

OpenStudy (anonymous):

and no matter what you do you are going to end up with \[(-\frac{b}{2a},f(-\frac{b}{2a}))\]

myininaya (myininaya):

\[y=ax^2+bx+c=a(x^2+\frac{b}{a}x)+c=a(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)+c-a(\frac{b}{2a})^2\] \[y=a(x+\frac{b}{2a})^2+c-\frac{b}{4a}\] \[y=a(x-(-\frac{b}{2a}))^2+(c-\frac{b}{4a})\] vertex is \[(-\frac{b}{2a},c-\frac{b}{4a})\]

OpenStudy (anonymous):

it is \[y=-\frac{b}{2a}\] how hard is that?

OpenStudy (anonymous):

So in a nutshell there are 3 ways to find the vertex. 1) \(\large x_{vertex} = -\frac{b}{2a}\) 2) Re-write in vertex form. 3) Find the midpoint between the 0's of the quadratic (if it has them)

OpenStudy (anonymous):

and pretty clearly method one is easiest

OpenStudy (anonymous):

Thanks so much.

OpenStudy (anonymous):

we all say you are welcome

OpenStudy (anonymous):

and we all like to fuss with each other. it is fun

myininaya (myininaya):

my is clearly the easiest to derive it each time

myininaya (myininaya):

lol

OpenStudy (anonymous):

oh yes. actually easiest is finding the derivative using the difference quotient. on an etchasketch

OpenStudy (anonymous):

easiest to do sure. easiest to remember, nope. The last one is easiest to remember (most intuitive), but not always reliable. The first one is easiest to do if you remember it. The middle one is a nice middle ground and always works.

OpenStudy (anonymous):

minus bee over two ay?

OpenStudy (anonymous):

man i thought my memory was bad...

OpenStudy (anonymous):

totally arbitrary data.

myininaya (myininaya):

like polpak said its hard too remember lol

OpenStudy (anonymous):

oh but it is not arbitrary at all. it is what you get

myininaya (myininaya):

satellite you just need to conform to my way

myininaya (myininaya):

putting it in vertex form is the bestest i win!

OpenStudy (anonymous):

\[y=ax^2+bx+c\] \[y=a(x+\frac{b}{2a})^2+\text{stuff}\] vertex is clearly when \[x=-\frac{b}{2a}\] because that makes the perfect square part zero

OpenStudy (anonymous):

I agree.

myininaya (myininaya):

+stuff?

OpenStudy (anonymous):

Vertex form is best.

OpenStudy (anonymous):

yeah stuff. i don't feel like keeping track. if i want "stuff" i replace x by \[-\frac{b}{2a}\] so see what i get

OpenStudy (anonymous):

what could be easier?

OpenStudy (anonymous):

See I already screwed it up in this coversation. I thought it was -a/2b

OpenStudy (anonymous):

who want to say "i added this but i really added that so i have to subtract the other"?

OpenStudy (anonymous):

Totally arbitrary.

myininaya (myininaya):

i want to say that

OpenStudy (anonymous):

You probably like foil too ;p

OpenStudy (anonymous):

@poplak look at this and tell me why it is arbitrary instead of obviously true \[y=ax^2+bx+c\] \[y=a(x+\frac{b}{2a})^2+\text{stuff}\]

OpenStudy (anonymous):

first, outer, inner, last, what could be easier?

OpenStudy (anonymous):

gimme a break there is a reason for it to be \[-\frac{b}{2a}\] foil is for turkeys at thanksgiving

OpenStudy (anonymous):

Because you're completing the square

OpenStudy (anonymous):

Which is basically solving for vertex form.

OpenStudy (anonymous):

Except you jump to the end and remember some formula -a/2b or some such.

myininaya (myininaya):

\[y=ax^2+bx+c=a(x^2+\frac{b}{a}x)+c=a(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)+c-a(\frac{b}{2a})^2\] \[y=a(x+\frac{b}{2a})^2+c-\frac{b}{4a} \] \[y=a(x-(-\frac{b}{2a}))^2+(c-\frac{b}{4a}) \] satellite why don't you just remember that the vertex is just (-\frac{b}{2a},c-\frac{b}{4a}) why do you only like to remember the x part only?

OpenStudy (anonymous):

You guys are still here...LOL

myininaya (myininaya):

\[(-\frac{b}{2a},c-\frac{b}{4a}) \]

OpenStudy (anonymous):

because i find it easy to evaluate a quadratic funciton

OpenStudy (anonymous):

NERD FIGHT!

OpenStudy (anonymous):

Commence flailing and hand slapping on the count of three!

OpenStudy (anonymous):

but at least you must agree that when you complete the square, not matter how many times you do it, the first coordinate will be \[-\frac{b}{2a}\] so there is nothing at all arbitrary about it

myininaya (myininaya):

but \[f(-\frac{b}{2a})=c-\frac{b}{4a}\] whats the point in pluggin' that if you are just gonna get what i got

OpenStudy (anonymous):

But you have to remember it that way instead of remembering 'complete the sqaure to get vertex form'. I can remember methods. I cannot remember formula.

OpenStudy (anonymous):

why would you want to repeat the procedure again and again. unless you are trying to teach someone how to do it

myininaya (myininaya):

why don't you remember everything satellite since remembering is so much easier

OpenStudy (anonymous):

Dyslexia is a feather I guess is what it boils down to.

OpenStudy (anonymous):

@polpak i do not actually believe you because if you wanted to solve \[3x^2+5x-1=0\] i am willing to bet that you will NOT complete the square

OpenStudy (anonymous):

frock together

OpenStudy (anonymous):

\[\text{feather} = \text{b}\text{i}\text{t}\text{c}\text{h}\]

myininaya (myininaya):

i would use my new method to factor that and i'm havent shown polpak my new method

OpenStudy (anonymous):

really? what language?

OpenStudy (anonymous):

In the silly anti-cursing filter they have here.

OpenStudy (anonymous):

assumption

OpenStudy (anonymous):

they seem to have fixed it

OpenStudy (anonymous):

pass the class

OpenStudy (anonymous):

don't be an retrice

OpenStudy (anonymous):

oh they fixed it

OpenStudy (anonymous):

apparently.

myininaya (myininaya):

what did they fix

OpenStudy (anonymous):

good because it was asinine

OpenStudy (anonymous):

In any event, I'm dyslexic. So I cannot remember random jumbles of letters and numbers in formula. I have to remember methods.

OpenStudy (anonymous):

oh you couldn't type any word with retricein it

OpenStudy (anonymous):

oh they really fixed it